OpenStudy (decarr432):

let f(x)=sqrt 7x and g(x)= x+8 what is the smallest number that is in the domain of (f o g)

OpenStudy (decarr432):

@pooja195 do you think you could help me?

OpenStudy (decarr432):

@DanJS

OpenStudy (astrophysics):

This is similar to your previous question, please set it up as f(g(x))

OpenStudy (decarr432):

Okay

OpenStudy (decarr432):

@Nnesha

OpenStudy (decarr432):

would that be sqrt(x+8)?

OpenStudy (astrophysics):

\[\huge f(g(x)) = \sqrt{7(x+8)}\]

OpenStudy (astrophysics):

Now how do you find the domain of this composition?

OpenStudy (decarr432):

Plug in numbers?

OpenStudy (astrophysics):

Well, we can't have a negative number in the square root right?

OpenStudy (decarr432):

No it can't be negative

OpenStudy (astrophysics):

Good, so simply we take what's inside our square root \[\huge 7(x+8) \ge 0\] solve for x :)

OpenStudy (decarr432):

but could my x for now be a negative to get say sqrt 7(-6+8) something like that and have 7(2) sqrt

OpenStudy (decarr432):

or would it not work

OpenStudy (astrophysics):

That does work, but plugging numbers would take forever, and if you solve for x, we can write our domain as it gives us a clear picture of what x's are allowed and what aren't. So go ahead and solve for x and you will see yourself!

OpenStudy (decarr432):

I have to get the smallest domain I can remember

OpenStudy (astrophysics):

That's ok, this will give you your domain. And we'll check after by plugging in numbers

OpenStudy (astrophysics):

To convince you

OpenStudy (decarr432):

Negative four would be the smallest tho

OpenStudy (astrophysics):

Please, solve for x

OpenStudy (decarr432):

How would you suggest I do that?

OpenStudy (astrophysics):

\[\huge 7(x+8) \ge 0 \] lets ignore the inequality sign for now, how would you solve for x if I gave you \[7(x+8)=0\]

OpenStudy (decarr432):

-8 right because

OpenStudy (decarr432):

\[7(-8+8)=0 7(0)=0\]

OpenStudy (astrophysics):

Exactly! So we have with our inequality \[x \ge - 8\]

OpenStudy (decarr432):

Okay now we have a starting point correct?

OpenStudy (astrophysics):

Well I would have hoped you do it algebraically, as that's what important, so to do it that way \[\frac{ 7(x+8)}{ 7 } \ge \frac{ 0 }{ 7 } \implies x+8 \ge 0 \implies x +8-8 \ge 0-8 \]\[\implies x \ge -8\]

OpenStudy (astrophysics):

Good?

OpenStudy (decarr432):

okay slow down slightly

OpenStudy (astrophysics):

Ok no worries, did you get lost anywhere during the algebraic process?

OpenStudy (decarr432):

Yes sir

OpenStudy (astrophysics):

Alright so our goal here was to solve for x (get x on its own) ok? We started off by dividing both sides by 7 as you can see the 7 is "attached" to the (x+8) meaning it is being multiplied, so we use the opposite operation which is division to remove it as it will turn it into 1 on the left side and 0/7 of course is 0.

OpenStudy (decarr432):

okay I get that part

OpenStudy (astrophysics):

Ok the next part we had \[x+8 \ge 0\] similarly notice how we are adding here on the left side, so again we use the opposite operation of what already exists and apply it to both sides. So we subtract -8 on both sides \[x+8 \color{red}{-8} \ge 0 \color{red}{-8} \implies x \ge \color{red}{-8}\] good?

OpenStudy (decarr432):

Yeah thanks for changing colors to show me btw

OpenStudy (astrophysics):

No worries, now let me ask you a question. What value is bigger -8 or -4?

OpenStudy (decarr432):

-4 is bigger

OpenStudy (decarr432):

Because its closer to zero

OpenStudy (astrophysics):

Right! :) Ok notice since we solved for x there, that inequality means x must be `greater than or equal` to -8, so that will be our smallest value allowed.

OpenStudy (decarr432):

okay but can the domain of f o g be a negative?

OpenStudy (decarr432):

because -8+8=0 Then 7(0)=0 sqrt(0)=0 would that work?

OpenStudy (astrophysics):

Yes, that is what we were solving for, don't confuse it with your last question because we had \[x-1 \ge 0 \implies x \ge 1\] in your last question

OpenStudy (astrophysics):

^ that's from your last question

OpenStudy (astrophysics):

If I can recall it correctly haha

OpenStudy (decarr432):

That is but he told me it has to be greater then 1

OpenStudy (astrophysics):

Ah gotcha, it was a denominator as well yes then it would be greater 1! Also, you are allowed to have 0 under your square root

OpenStudy (decarr432):

Okay thank you for clearing that up

OpenStudy (astrophysics):

Yw :)

OpenStudy (decarr432):

So my smallest number would be -8 correct because the square of 0 is 0

OpenStudy (astrophysics):

Correct, essentially when you have a square root, we take what's inside it and set it up an inequality as we did before and that will give you your restrictions for your domain.

OpenStudy (decarr432):

okay thanks

OpenStudy (astrophysics):

:)

OpenStudy (decarr432):

The answer was correct thanks a lot :D

OpenStudy (decarr432):

now I have a tough question lol